OpenStudy (anonymous):

The length of a rectangle is eight more than thrice the width. Write an expression for the area of the rectangle.

OpenStudy (kkutie7):

A=L*W

OpenStudy (kkutie7):

A=w*3w+8

OpenStudy (anonymous):

L=8+3*W A=L*(8+3*W)

OpenStudy (anonymous):

Thank You..... <3